the student council of sta lucia high school needed 2250.00 for an annual trip. If five more members joined the group, each member would then save 5.00. How many students planned to make the trip?
Answer by lwsshak3(11628)

the student council of sta lucia high school needed 2250.00 for an annual trip. If five more members joined the group, each member would then save 5.00. How many students planned to make the trip?
**
let x=no of members who planned to make the trip
Initial member cost of trip=2250/x
member cost of trip with 5 more members joining=2250/(x+5)
2250/x-2250/(x+5)=5
LCD: x(x+5)
2250(x+5)+2250x=5x(x+5)
2250x+5*2250-2250x=5x^2+25x
5*2250=5x^2+25x
2250=x^2+5x
x^2+5x-2250=0
(x+50)(x-45)=0
x=-50 (reject, x>0)
or
x=45
no. of members who planned to make the trip=45